# Index Terms: V2V CSI, Measurements, Spatiotemporal Predictive Learning, Context-Aware Attention, and Pseudo-Labeling Optimization.I Introduction

### v1

@@ Abstract
Vehicle-to-vehicle (V2V) communications are crucial for future driving, especially for assisted or autonomous systems [1, 2, 3, 4]. These systems enable vehicles to warn each other of imminent actions like emergency braking or coordinate smooth lane changes. However, widespread adoption of V2V communication has been slow due to, at least partially, economic factors and the unpredictable performance of V2V systems. The latter can be attributed to challenges that include signal propagation issues and high device density, causing interference and packet loss [5, 6]. Thus, improving the reliability and latency of V2V links is essential. Due to the high dynamics in V2V, a key challenge is to maintain a robust communication with outdated channel measurement. Therefore, effective channel prediction methods are needed to infer the current state from past data.
@@ II-A Machine Learning based V2V Channel Prediction
II-B On the Predictive Learning Algorithms
@@
In summary, each type of method has its own strengths and limitations. Transformers are highly effective at extracting semantic correlations in long sequences; However, in multiple dimensional time series modeling, the goal is to capture temporal relationships within an ordered sequence of continuous points in spatial-temporal domains. Although positional encoding and token embeddings help maintain some ordering, the permutation-invariant self-attention mechanism inevitably leads to a loss of temporal information, as demonstrated in [42]. On the other hand, ConvLSTM and its variants are highly effective at modeling both spatial and temporal data, demonstrating strong performance in spatiotemporal prediction tasks. Nevertheless, they can be susceptible to the accumulated prediction error(APE)[43, 44].
@@ II-C Our Contributions
With the motivations mentioned above, we propose a novel predictive learning method for realistic V2V channel prediction, focusing on the built-in properties of V2V data and leveraging well-established spatio-temporal predictive learning models. Our key contributions are summarized as follows:
@@ III-B Challenges in Multi-Dimensional V2V Channel Prediction
III-B1 The Built-in Properties of the V2V Channel Measurements
This subsection details the data structure of the V2V CSI measurements collected using our channel sounder, initially introduced in [45]. The transmitter (Tx) and receiver (Rx) each consist of 8-element vertically polarized uniform circular dipole arrays mounted on vehicles. During the measurement campaigns, detailed further in Section V-A, the Tx and Rx communicate over a system with a 5.9 GHz carrier frequency and a 15 MHz bandwidth. The maximum resolvable Doppler shift is given by , approximately 806 Hz, which corresponds to a maximum relative speed of around 148 km/h. We measure the related MIMO channel burst by burst, each burst containing 30 snapshots, where one snapshot captures the complete MIMO channel, i.e., the transfer function between each Tx and Rx element. Bursts are repeated every 50 ms. The MIMO sounding signal comprises 64 (8 × 8) repetitions of this signal, with a total duration of 640 s. As explained in [45], our setup consists of a pair of NI-USRP RIOs serving as the main RF transceivers, along with a pair of 8-element switched antenna arrays. Several guard periods are inserted between the sounding signals to accommodate the settling time of the Tx or Rx switches.
@@ IV The Proposed Method
Building on these motivations, this section details the proposed approach for addressing the complex problem of multi-dimensional CSI prediction. The overall framework is given in Fig. 1. We begin by introducing our predictive learning model, which is tailored to capture dependencies in multiple dimensional CSI sequences effectively. Furthermore, we introduce the network optimization method, which is based on the meta learning framework, incorporating the concept of meta pseudo labels to enhance network training and mitigate the APE bottleneck issue.
@@ IV-A Preliminaries
For a better understanding of the proposed method, we first explain the CA-ConvLSTM module, which includes two core parts: The casual ConvLSTM and the gradient highway. A CA-ConvLSTM unit features dual memories: the temporal memory and the spatial memory . In this notation, the subscript represents the time step, while the superscript denotes the -th hidden layer. The current temporal memory directly depends on its previous state and is regulated through three gates: a forget gate , an input gate , and an input modulation gate . Meanwhile, the current spatial memory is influenced by in the deeper transition path. Specifically, for the bottom layer (), the topmost spatial memory at is assigned to . Distinct from the original spatiotemporal LSTM, the CA-ConvLSTM utilizes a cascaded mechanism, where the spatial memory is particularly a function of the temporal memory through an additional set of gate structures. The update equations of the CA-ConvLSTM at the -th layer are as follows:
@@ IV-C3 A Minor Refinement: the Proposed Adaptive Teacher
The above optimization, as shown in Section IV-C2, is based on MSE loss (). Here we provide a minor refinement for it ass proposing a weighted MSE loss function, which is given by
@@ V-A1 A brief introduction to the measurement campaign
In our measurement campaigns, we utilized a channel sounder, as described in [45], to measure the CSI across three challenging scenarios: Scenario I: Mixed City and Campus Road - This scenario combines urban and campus road environments; Scenario II: Campus Road - This scenario focuses solely on the campus road environment; Scenario III: Highway - This scenario pertains to a highway setting. The trajectories of the transmitter (TX) and receiver (RX) for each of these scenarios are depicted in Fig. 3. In Scenarios I and III, both TX and RX were mobile, moving along predefined paths. Conversely, in Scenario II, the RX remained stationary while the TX was in motion. These configurations were intentionally designed to explore the different mobility patterns in Vehicle-to-Vehicle (V2V) communication systems. A comprehensive summary of the measured CSI data for each scenario is provided in Tab. II.
@@ V-A2 Performance Analysis
For our comparison of spatio-temporal predictive learning algorithms, we focus on evaluating the proposed method against two prominent algorithms: ConvLSTM [38], and ST-ConvLSTM [54]. ConvLSTM is a well-established algorithm, extensively cited (over 9,400 citations), and has been validated for its effectiveness in CSI prediction problems within wireless communication systems [23]. It utilizes convolutional structures within LSTM units to capture both spatial and temporal dependencies, making it particularly suitable for our evaluation. ST-ConvLSTM, on the other hand, represents the state-of-the-art in spatio-temporal predictive learning algorithms. ST-ConvLSTM employs an advanced recurrent neural network with novel spatio-temporal memory designs specifically designed to handle complex spatio-temporal dynamics, offering state-of-the-art predictive performance over numerous methods, as verified in [33]. By comparing our proposed method with these two advanced algorithms, we aim to provide a comprehensive assessment of its effectiveness in the context of spatio-temporal predictive learning for CSI prediction.
@@ V-B Prediction Performance Over Time
For the production results related to Scenario I (City and campus roads), shown in the top left and bottom left graphs, the proposed method slightly outperforms ST-ConvLSTM and ConvLSTM, maintaining lower error values across the timesteps. In this low-mobility scenario, both the transmitter (TX) and receiver (RX) were moving along the USC campus road. Due to the large coherence time, the channel characteristics remain relatively stable [29]. As a result, all the predictive learning algorithms exhibit only slight variations in their prediction performances, reflecting the minimal impact of mobility on the channel conditions in this specific context.
In Scenario II (Campus Roads), the channel undergoes significant changes due to the differing mobility of the TX and RX. The corresponding prediction results indicate that the proposed method maintains superior performance with lower error values. However, the margin of improvement is more significant than in Scenario I. This underscores the effectiveness of the advanced design of context-conditioned memory updates in spatiotemporal learning algorithms, confirming their capability to handle varying mobility conditions.
@@ V-C1 CSI Prediction With Supervised Learning
As a comparison, we also present the comprehensive prediction results in a cross-geo setting, where all predictive learning models (ST-ConvLSTM and the proposed method) are trained on data from one scenario but tested on data from a different scenario. Figure 6 shows that all predictive learning models experience significant performance loss in this challenging setting. This performance degradation is primarily due to the inherent difficulties of V2V channel prediction with measurements collected from different scenarios. Applying trained models from one scenario to data from another proves challenging due to the unique characteristics of CSI data under varying propagation conditions, leading to a severe APE issue. Moreover, we observe that mobility patterns significantly impact prediction performance. For instance, in scenarios with similar moving patterns (S1 and S3), the results for S1S3 and S3S1 (Fig. 6a and Fig. 6c) are better than those for S2S3 and S2S1 (Fig. 6b). This suggests that spatiotemporal predictive learning models can effectively capture the moving patterns, although it remains challenging to apply the learned patterns from one scenario to another. For more accurate and reliable results, we will present the outcomes in the meta learning setting below.
@@ V-C2 CSI Prediction With meta learning
We present the prediction results using a cross-geo setup, where the network is trained with labeled data from one scenario and tested with data from a different (unseen) scenario. For the meta learning setup, some unlabeled data from the unseen scenario is also included in the training process. Fig. 7a corresponds to the scenario S1-S2, while Fig. 7b corresponds to S1-S3. We use the CDF curves of MSEs for the proposed method within several learning frameworks, including ”Without meta learning learning”, ”With meta learning”, and ”With adaptive meta learning learning”, which have been introduced in Section IV-C. For the ”Supervised” setup, we consider the fully supervised results (same-geo) as a performance benchmark for the meta learning methods.
As shown in Fig.7, the proposed method with meta learning schemes, such as ”With meta learning” and ”With adaptive meta learning,” demonstrates superior performance, with ”With adaptive meta learning” showing slightly better results due to its adaptive nature. Although there is a performance loss compared to the fully supervised benchmark, this loss is less than the loss between ”With meta learning” and ”Without meta learning” case. These findings reveal that, in a cross-geo setup, there is considerable potential for performance improvement in predictive learning algorithms, primarily due to the APE bottleneck issue. meta learning schemes can effectively address this issue by generating and optimizing pseudo labels for unlabeled data, significantly reducing the APE. In summary, the case study in Fig. 7 underscores the importance of leveraging meta learning techniques to enhance prediction accuracy in complex and dynamic scenarios. Additional details, such as the applicability of meta learning for different predictive models and the impact of the number of available labeled samples, will be provided in the following section.
@@ VI Conclusions
In this paper, we introduced a novel context-conditioned spatiotemporal predictive model for the challenging V2V channel prediction problem using measurement data. We proved the effectiveness of context-conditioned attention, which considers the built-in properties of CSI data, by presenting comprehensive prediction results across various scenarios, ranging from low to high mobility. Moreover, we explained the bottleneck APE issue in recurrent-based predictive models and demonstrated the effectiveness of a meta learning framework in addressing this issue. We verified that meta learning can be applied to all predictive models, albeit at the cost of a higher training budget.

### v2

@@ Abstract
Vehicle-to-vehicle (V2V) communications are crucial for future driving, especially for assisted or autonomous systems [1, 2, 3, 4]. These systems enable vehicles to warn each other of imminent actions like emergency braking or coordinate smooth lane changes. However, widespread adoption of V2V communication has been slow due to, at least partially, economic factors and the unpredictable performance of V2V systems. The latter can be attributed to challenges that include signal propagation issues and high device density, causing interference and packet loss [5, 6]. Thus, improving the reliability and latency of V2V links is essential. Due to the high dynamics in V2V environments, a key challenge is to maintain a robust communication with outdated channel measurements. Therefore, effective channel prediction methods are needed to infer the current state from past data.
@@ II-A Machine Learning based V2V Channel Prediction
II-B Predictive Learning Algorithms
@@
In summary, each type of method has its own strengths and limitations. Transformers are highly effective at extracting semantic correlations in long sequences; However, in multiple dimensional time series modeling, the goal is to capture temporal relationships within an ordered sequence of continuous points in spatial-temporal domains. Although positional encoding and token embeddings help maintain some ordering, the permutation-invariant self-attention mechanism may lead to a loss of temporal information, as demonstrated in [42]. On the other hand, ConvLSTM and its variants are highly effective at modeling both spatial and temporal data, demonstrating strong performance in spatiotemporal prediction tasks. Nevertheless, they can be susceptible to the accumulated prediction error (APE)[43, 44].
@@ II-C Our Contributions
With the motivations mentioned above, we propose a novel predictive learning method for realistic V2V channel prediction, focusing on the inherent properties of V2V data and leveraging well-established spatio-temporal predictive learning models. Our key contributions are summarized as follows:
@@ III-B Challenges in Multi-Dimensional V2V Channel Prediction
III-B1 The Inherent Properties of the V2V Channel Measurements
This subsection details the data structure of the V2V CSI measurements collected using our channel sounder, initially introduced in [45]. The transmitter (Tx) and receiver (Rx) each consist of 8-element vertically polarized uniform circular dipole arrays mounted on vehicles. During the measurement campaigns, detailed further in Section V-A, the Tx and Rx communicate over a system with a 5.9 GHz carrier frequency and a 15 MHz bandwidth. As explained in [45], our setup consists of a pair of NI-USRP RIOs serving as the main RF transceivers, along with a pair of 8-element switched antenna arrays. The MIMO sounding signal comprises 64 (8 × 8) repetitions of a multi-tone sounding signal, with a total duration of 640 s. Several guard periods are inserted between the sounding signals to accommodate the settling time of the Tx or Rx switches. The maximum resolvable Doppler shift is given by , approximately 806 Hz, which corresponds to a maximum relative speed of around 148 km/h. We measure the related MIMO channel burst by burst, each burst containing 30 snapshots, where one snapshot captures the complete MIMO channel, i.e., the transfer function between each Tx and Rx element. Bursts are repeated every 50 ms.
@@ IV The Proposed Method
Building on these motivations, this section details the proposed approach for addressing the complex problem of multi-dimensional CSI prediction. The overall framework is given in Fig. 1. We begin by introducing our predictive learning model, which is tailored to capture dependencies in multi-dimensional CSI sequences effectively. Furthermore, we introduce the network optimization method, which is based on the meta learning framework, incorporating the concept of meta pseudo labels to enhance network training and mitigate the APE bottleneck issue.
@@ IV-A Preliminaries
For a better understanding of the proposed method, we first explain the CA-ConvLSTM module, which includes two core parts: The causal ConvLSTM and the gradient highway. A CA-ConvLSTM unit features dual memories: the temporal memory and the spatial memory . In this notation, the subscript represents the time step, while the superscript denotes the -th hidden layer. The current temporal memory directly depends on its previous state and is regulated through three gates: a forget gate , an input gate , and an input modulation gate . Meanwhile, the current spatial memory is influenced by in the deeper transition path. Specifically, for the bottom layer (), the topmost spatial memory at is assigned to . Distinct from the original spatiotemporal LSTM, the CA-ConvLSTM utilizes a cascaded mechanism, where the spatial memory is particularly a function of the temporal memory through an additional set of gate structures. The update equations of the CA-ConvLSTM at the -th layer are as follows:
@@ IV-C3 A Minor Refinement: the Proposed Adaptive Teacher
The above optimization, as shown in Section IV-C2, is based on MSE loss (). Here we provide a minor refinement for it by proposing a weighted MSE loss function, which is given by
@@ V-A1 A brief introduction to the measurement campaign
In our measurement campaigns, we utilized a channel sounder, as described in Sec. III.B and in more detail in [45], to measure the CSI across three challenging scenarios: Scenario I: Mixed City and Campus Road - This scenario combines urban and campus road environments; Scenario II: Campus Road - This scenario focuses solely on the campus road environment; Scenario III: Highway - This scenario pertains to a highway setting. The trajectories of the transmitter (TX) and receiver (RX) for each of these scenarios are depicted in Fig. 3. In Scenarios I and III, both TX and RX were mobile, moving along predefined paths. Conversely, in Scenario II, the RX remained stationary while the TX was in motion. These configurations were intentionally designed to explore the different mobility patterns in Vehicle-to-Vehicle (V2V) communication systems. A summary of the measured CSI data for each scenario is provided in Tab. II.
@@ V-A2 Performance Analysis
For our comparison of spatio-temporal predictive learning algorithms, we focus on evaluating the proposed method against two prominent algorithms: ConvLSTM [38], and ST-ConvLSTM [54]. ConvLSTM is a well-established algorithm, extensively cited, and has been validated for its effectiveness in CSI prediction problems within wireless communication systems [23]. It utilizes convolutional structures within LSTM units to capture both spatial and temporal dependencies, making it particularly suitable for our evaluation. ST-ConvLSTM, on the other hand, represents the state-of-the-art in spatio-temporal predictive learning algorithms. ST-ConvLSTM employs an advanced recurrent neural network with novel spatio-temporal memory designs specifically designed to handle complex spatio-temporal dynamics, offering state-of-the-art predictive performance over numerous methods, as verified in [33]. By comparing our proposed method with these two advanced algorithms, we aim to provide a comprehensive assessment of its effectiveness in the context of spatio-temporal predictive learning for CSI prediction.
@@ V-B Prediction Performance Over Time
For the results related to Scenario I (City and campus roads), shown in the top left and bottom left graphs, the proposed method slightly outperforms ST-ConvLSTM and ConvLSTM, maintaining lower error values across the timesteps. In this low-mobility scenario, both the TX and RX were moving along the USC campus road. Due to the large coherence time, the channel characteristics remain relatively stable [29]. As a result, all the predictive learning algorithms exhibit only slight variations in their prediction performances, reflecting the minimal impact of mobility on the channel conditions in this specific context.
In Scenario II (Campus Roads), the channel undergoes significant changes due to the differing mobility of the TX and RX. The corresponding prediction results indicate that the proposed method maintains superior performance with lower error values; the margin of improvement is more significant than in Scenario I. This underscores the effectiveness of the advanced design of context-conditioned memory updates in spatiotemporal learning algorithms, confirming their capability to handle varying mobility conditions.
@@ V-C1 CSI Prediction With Supervised Learning
As a comparison, we also present the comprehensive prediction results in a cross-geo setting, where all predictive learning models (ST-ConvLSTM and the proposed method) are trained on data from one scenario but tested on data from a different scenario. Figure 6 shows that all predictive learning models experience significant performance loss in this challenging setting. This performance degradation is primarily due to the inherent difficulties of V2V channel prediction with measurements collected from different scenarios. Applying trained models from one scenario to data from another proves challenging due to the unique characteristics of CSI data under varying propagation conditions, leading to a severe APE issue. Moreover, we observe that mobility patterns significantly impact prediction performance. For instance, cross-scenario training with similar moving patterns, i.e., S1S3 and S3S1 (Fig. 6a and Fig. 6c), provides better results than those for different mobility patterns, S2S3 and S2S1 (Fig. 6b). This suggests that spatiotemporal predictive learning models can effectively capture the moving patterns, although it remains challenging to apply the learned patterns from one scenario to another. For more accurate and reliable results, we will present the outcomes in the meta learning setting below.
@@ V-C2 CSI Prediction With meta learning
We present the prediction results using a cross-geo setup, where the network is trained with labeled data from one scenario and tested with data from a different (unseen) scenario. For the meta learning setup, some unlabeled data from the unseen scenario are also included in the training process. Fig. 7a corresponds to the scenario S1-S2, while Fig. 7b corresponds to S1-S3. We use the CDF curves of MSEs for the proposed method within several learning frameworks, including ”Without meta learning learning”, ”With meta learning”, and ”With adaptive meta learning learning”, which have been introduced in Section IV-C. For the ”Supervised” setup, we consider the fully supervised results (same-geo) as a performance benchmark for the meta learning methods.
As shown in Fig.7, the proposed methods with meta learning schemes, such as ”With meta learning” and ”With adaptive meta learning,” demonstrates superior performance, with ”With adaptive meta learning” showing slightly better results due to its adaptive nature. Although there is a performance loss compared to the fully supervised benchmark, this loss is less than the loss between ”With meta learning” and ”Without meta learning” case. These findings reveal that, in a cross-geo setup, there is considerable potential for performance improvement in predictive learning algorithms, primarily due to the APE bottleneck issue. Meta learning schemes can effectively address this issue by generating and optimizing pseudo labels for unlabeled data, significantly reducing the APE. In summary, the case study in Fig. 7 underscores the importance of leveraging meta learning techniques to enhance prediction accuracy in complex and dynamic scenarios. Additional details, such as the applicability of meta learning for different predictive models and the impact of the number of available labeled samples, will be provided in the following section.
@@ VI Conclusions
In this paper, we introduced a novel context-conditioned spatiotemporal predictive model for the challenging V2V channel prediction problem using measurement data. We proved the effectiveness of context-conditioned attention, which considers the inherent properties of CSI data, by presenting comprehensive prediction results across various scenarios, ranging from low to high mobility. Moreover, we explained the bottleneck APE issue in recurrent-based predictive models and demonstrated the effectiveness of a meta learning framework in addressing this issue. We verified that meta learning can be applied to all predictive models, albeit at the cost of a higher training budget.
@@
Acknowledgment
The authors thank the proofreading from Omer Serbetci and the valuable discussions with members of the WiDeS group, including Dr. Naveed Abbasi, Dr. Yang Cai, Weiyu Chen, Zihang Cheng, Dr. Hussein Hammoud, Dr. Juhyung Lee, and Dr. Jorge Gomez Ponce. The authors appreciate the support from Dr. Rui Wang in channel sounder and data prepossessing.